OpenStudy (11study11):

Just check answers I'm desperate!!!!!!! please!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Medals + fan

OpenStudy (11study11):

Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 4^–x and y = 2x + 3 intersect are the solutions of the equation 4–x = 2x + 3. (4 points) Part B: Make tables to find the solution to 4^–x = 2x + 3. Take the integer values of x between −3 and 3. (4 points) Part C: How can you solve the equation 4^–x = 2x + 3 graphically? (2 points)

OpenStudy (11study11):

Part A: Where the two lines intersect is the answer to both solutions because it is a point for both y = 4^–x and y = 2x + 3. Part B: x y = 4^–x y = 2x + 3 -3 y= 64 y=-3 -2 y=16 y=-1 -1 y=4 y=1 0 y= 1 y=3 1 y=0.25 y=5 2 y=0.0625 y=7 3 y=0.015625 y=9 Part C: You would draw the lines on the graph and the solution will be the point where the two lines intersect.

OpenStudy (11study11):

@HolsterEmission @Google

OpenStudy (11study11):

@Arcadiouse @thor123 @brian22

OpenStudy (11study11):

please help I'm desperate!

jhonyy9 (jhonyy9):

@11study11 you are right with this answers sure hope helped